# Denoising Diffusion Variational Inference: Diffusion Models as Expressive Variational Posteriors

### v3

@@ 3.2 The Wake-Sleep Regularized ELBO
The standard approach to fit auxiliary-variable generative models [19] is to apply the ELBO twice:
$\log p_{\bm{\theta}}(\mathbf{x})$ $\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))$ (5)
$\geq\log p_{\bm{\theta}}(\mathbf{x})-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))-\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$ (6)
$=~{}\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$ (7)
@@
In Equation 5, we applied the ELBO over $\mathbf{z}$, and in Equation 6 we applied the ELBO again over the latent $\mathbf{y}$ of $q$. Notice that the gap between the ELBO and $\log p_{\bm{\theta}}(\mathbf{x})$ is $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})||p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}[D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$. Thus, if we correctly match $q$ and $r$, we will achieve a tight bound.
Wake-Sleep Regularization
Notice that optimizing $D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{z},\mathbf{x})p(\mathbf{z}))$ in Equation 7 involves sampling from the approximate reverse process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ to match the true reverse process $r(\mathbf{y}|\mathbf{z},\mathbf{x})$: this is the opposite of diffusion training, where we would sample from $r$ to fit $q$. We have found the standard ELBO to insufficient to learn a good $q_{\bm{\phi}}$ that reverses the noising process $r$, as illustrated in Table 8.
Instead, our strategy is to introduce a modified ELBO that yields diffusion-like training. This objective is the ELBO in Equation 7 augmented with an additional regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$.
$\log p_{\bm{\theta}}(\mathbf{x})\geq\underbrace{\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})]}_{\text{wake / recons. term $\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))}_{\text{prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$}}\underbrace{-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]}_{\text{sleep term $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$}}$ (8)
The optimization of the regularizer $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ is similar to the sleep phase of the wake-sleep algorithm, and closely resembles diffusion model training (see below). As in wake-sleep, $\mathcal{L}_{\text{sleep}}$ is optimized over ${\bm{\phi}}$ only, and the $\mathbf{x}$ are sampled from the model.
3.3 Optimizing the Regularized ELBO
Computing $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ still involves intractable distributions $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x}),q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$. To optimize $\mathcal{L}_{\text{sleep}}({\bm{\phi}})$, we introduce another lower bound $\mathcal{L}_{\text{diff}}({\bm{\phi}})$, which we call the denoising diffusion loss (for reasons that will become apparent shortly):
$\mathcal{L}_{\text{sleep}}({\bm{\phi}})$ $=-\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x})}[D_{\mathrm{KL}}(p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})||q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}))]=\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})$ (9)
$\geq\mathbb{E}_{p_{\bm{\theta}}(\mathbf{x},\mathbf{z})}[\mathbb{E}_{r}[\log\frac{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}{r(\mathbf{y}|\mathbf{z},\mathbf{x})}]+\bar{H}(p_{\bm{\theta}})=\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (10)
In Equation 10, we applied the ELBO with $r(\mathbf{y}|\mathbf{z},\mathbf{x})$ playing the role of the variational posterior over the latent $\mathbf{y}$ in $q_{\bm{\phi}}$; $\bar{H}(p_{\bm{\theta}})$ is the expected conditional entropy of $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, a constant that does not depend on ${\bm{\phi}}$.
We define the DDVI learning objective $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ as the sum of the aforementioned terms:
$\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})=\mathcal{L}_{\text{rec}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{reg}}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})+\mathcal{L}_{\text{diff}}({\bm{\phi}})$ (11)
Terms $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{diff}}$ may be weighted by hyper-parameters $\beta_{\text{reg}},\beta_{\text{diff}}>0$, as in the $\beta$-VAE framework. In our experiments, assume $\beta_{\text{reg}}=\beta_{\text{diff}}=1$ unless otherwise specified. Note that since $\mathcal{L}_{\text{diff}}\leq\mathcal{L}_{\text{sleep}}\leq 0$, $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ is a valid lower bound on $\log p_{\bm{\theta}}(\mathbf{x})$ that is tight when $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})=p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$.
Wake-Sleep Optimization
We optimize our bound on $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ using gradient descent by alternating between ELBO optimization and taking sleep steps (see Section 5). Note that by maximizing $\mathcal{L}_{\text{diff}}$, we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the forward KL divergence; similarly, by optimizing $\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{reg}}$ (the ELBO), we fit $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$ via the reverse KL divergence. Thus, optimizing $\mathcal{L}(\mathbf{x},{\bm{\theta}},{\bm{\phi}})$ encourages $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ to approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, and when the two are equal, the bound on $\log p_{\bm{\theta}}(\mathbf{x})$ is tight.
Simplifying Wake-Sleep
We also consider a light-weight algorithm, in which $r(\mathbf{y}|\mathbf{z})$ and $q_{\bm{\phi}}(\mathbf{z}|\mathbf{y})$ do not depend on $\mathbf{x}$. In this case, $\mathcal{L}_{\text{diff}}$ requires only sampling from $p(\mathbf{z})$, and the entire loss $\mathcal{L}$ can be optimized end-to-end using gradient descent. This algorithm is a simpler (there is no separate sleep phase); however, $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})$ may not perfectly approximate $p_{\bm{\theta}}(\mathbf{z}|\mathbf{x})$, hence $\mathcal{L}$ may no longer be a tight bound. We report results in Appendix B.
3.4 Combining The Wake-Sleep Regularized ELBO With Diffusion Models
Lastly, we can further simplify $\mathcal{L}_{\text{diff}}$ by leveraging the Markov structure of the forward and reverse processes $r,q$. Recall that each $\mathbf{y}=(\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{T})$ can be a vector of $T$ latents, which we also denote as $\mathbf{y}_{1:T}$, and that $r(\mathbf{y}_{1:T}|\mathbf{z})=\prod_{t=1}^{T}r(\mathbf{y}_{t}|\mathbf{y}_{t-1})$, where $\mathbf{y}_{0}=\mathbf{z}$ and the $\mathbf{y}_{1:T}$ are increasingly noised versions of $\mathbf{y}_{0}$. Similarly, $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})$ is an approximate reverse diffusion process $q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{0:T}|\mathbf{x})=q_{\bm{\phi}}(\mathbf{y}_{T}|\mathbf{x})\prod_{t=1}^{T}q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ that is trained to reverse $r$.
In order to fit $q_{\bm{\phi}}$, we form a lower bound $\mathcal{L}_{\text{diff}}(\mathbf{x},{\bm{\phi}})$ on the sleep term $\mathbb{E}_{p(\mathbf{x},\mathbf{z})}\log{q_{\bm{\phi}}(\mathbf{z}|\mathbf{x})}$ in Equation 9. This bound is identical to the ELBO of a diffusion model, and has the same derivation.
$\mathcal{L}_{\text{diff}}=$ $~{}\mathbb{E}_{r}\left[\log q_{\bm{\phi}}(\mathbf{z}|\mathbf{y}_{1},\mathbf{x})-\sum_{t=2}^{T}D_{\mathrm{KL}}(r_{t}||q_{t})\right]-D_{\mathrm{KL}}(r(\mathbf{y}_{T}|\mathbf{z})||q_{\bm{\phi}}(\mathbf{y}|\mathbf{x})).$ (12)
where $r_{t},q_{t}$ denote the distributions $r(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{y}_{0})$ and $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$.
Parameterizing Diffusion-Based Encoders
A common type of noising process compatible with this bound when $\mathbf{z}$ is continuous is Gaussian diffusion, where we define $r(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{1-{\bm{\alpha}}_{t}}\mathbf{y}_{t-1},{\bm{\alpha}}_{t}\mathbf{I})$ for a suitable schedule $({\bm{\alpha}}_{t})_{t=1}^{T}$. We then adopt the parameterization $q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})=\mathcal{N}(\mathbf{y}_{t-1};\mu_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t),\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x},t))$. It is then common to parameterize $q_{\bm{\phi}}$ with a noise prediction network ${\bm{\epsilon}}_{\bm{\phi}}$ [12]; the sum of KL divergences can be approximated by $\mathbb{E}_{t,{\bm{\epsilon}}_{t}\sim r(\mathbf{y}_{0},t)}||{\bm{\epsilon}}_{t}-{\bm{\epsilon}}_{\bm{\phi}}(\sqrt{\bar{{\bm{\alpha}}_{t}}}\mathbf{y}_{0}+\sqrt{1-\bar{{\bm{\alpha}}_{t}}}{\bm{\epsilon}}_{t},\mathbf{x},t)||^{2}$.
Lastly, in order to use diffusion-based encoders with our ELBO-based objective, we need to show that we can tractably compute $-D_{\mathrm{KL}}(q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$, which we equivalently rewrite as:
$-D_{\mathrm{KL}}(q_{\bm{\phi}}||r\cdot p)=\mathbb{E}_{q_{\bm{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log(r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))]+H(q).$
The term $H(q)$ denotes the entropy. We approximate the first term using Monte-Carlo; the entropy is computed as
$H(q)=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q_{\bm{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]=\sum_{t=1}^{T+1}\mathbb{E}_{q}\left[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log\left|\Sigma_{\bm{\phi}}(\mathbf{y}_{t},\mathbf{x})\right|\right]$ (13)

### v4

@@ 3.2 Learning Objective: A Markovian ELBO
The standard approach to fit auxiliary-variable generative models (Maaløe et al. 2016) is to apply the ELBO twice:
$\log p_{\boldsymbol{\theta}}(\mathbf{x})$ $\geq\log p_{\boldsymbol{\theta}}(\mathbf{x})-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$ (3)
$\geq\log p_{\boldsymbol{\theta}}(\mathbf{x})-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))$ (4)
$~{}~{}~{}~{}-\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$
$=~{}\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})]$ (5)
$~{}~{}~{}~{}-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$
@@
In Equation (3), we applied the ELBO over $\mathbf{z}$, and in Equation (4) we applied the ELBO over the latent $\mathbf{y}$ of $q$ (see Appendix J for the derivation). Notice that the gap between the ELBO and $\log p_{\boldsymbol{\theta}}(\mathbf{x})$ is $\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})||p_{\boldsymbol{\theta}}(\mathbf{z}|\mathbf{x}))+\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x})}[\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y}|\mathbf{x},\mathbf{z})||r(\mathbf{y}|\mathbf{x},\mathbf{z}))]$. Thus, if we correctly match $q$ and $r$, we will achieve a tight bound.
Analyzing the ELBO
Optimizing Equation (5) requires tractably dealing with the prior regularization term $\mathcal{L}_{\text{reg}}(\mathbf{x},{\boldsymbol{\theta}},{\boldsymbol{\phi}}):=-\text{KL}(q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})||r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))$, which we equivalently rewrite as:
$\mathcal{L}_{\text{reg}}=\mathbb{E}_{q_{\boldsymbol{\phi}}(\mathbf{y},\mathbf{z}|\mathbf{x})}[\log(r(\mathbf{y}|\mathbf{x},\mathbf{z})p(\mathbf{z}))]+H(q).$ (6)
We can expand the first term by leveraging the Markov structure of $r,q$ to rewrite $\mathcal{L}_{\text{reg}}$ as the likelihood of samples from the reverse diffusion process $q$ under the forward process $r$.
$\mathcal{L}_{\text{reg}}=\sum_{t=1}^{T}\mathbb{E}_{q}[\log(r(\mathbf{y}_{t}|\mathbf{y}_{t-1},\mathbf{x})]+\mathbb{E}_{q}[\log p(\mathbf{z})]+H(q),$
where $\mathbf{y}_{0}:=\mathbf{z}$. We refer to optimizing the Markovian ELBO as unregularized DDVI, the first instance of our method. The noise process $r$ defines prior regularization terms for each $\mathbf{y}_{t}$. This provides extra supervision for learning $q$ in the form of trajectories from latents $\mathbf{y}_{T}$ to $\mathbf{y}_{1}$; this extra supervision helps $q$ fit complex non-Gaussian posteriors.
The term $H(q)=-\sum_{t=1}^{T+1}\mathbb{E}_{q}[\log q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})]$ denotes the entropy. For example, when each term $q_{\boldsymbol{\phi}}(\mathbf{y}_{t-1}|\mathbf{y}_{t},\mathbf{x})$ is Gaussian, it is computed as
$H(q)=$ $\sum_{t=1}^{T+1}\mathbb{E}_{q}\left[\frac{d}{2}\left(1+\log(2\pi)\right)+\frac{1}{2}\log\left|\Sigma_{\boldsymbol{\phi}}(\mathbf{y}_{t},\mathbf{x})\right|\right]$